# DEEM: Dynamic Experienced Expert Modeling for Stance Detection

### v1

@@ 1 Introduction
We evaluate DEEM across both single-target and multi-target stance detection tasks on three widely used datasets, including P-Stance (Li et al., 2021), SemEval-2016 (Mohammad et al., 2016), and MTSD (Sobhani et al., 2017). Experimental results demonstrate that DEEM with dynamic experienced experts can gain substantial improvement across all datasets. Furthermore, it also outperforms reasoning with self-consistency that requires multiple responses and shows potential for reducing the bias of LLMs.
@@ 3 Methods
[…]
@@ 3.1 Generating Diverse Experts
$\mathcal{E}_{j},\hat{l}_{j}={\mathcal{M}}({\rm prompt}\oplus s_{j}\oplus t_{j}),M$ (2)
@@ 4.1 Experimental Setups
FT BiCond (Augenstein et al., ${}^{\spadesuit}$ 73.0 69.4 64.6  ${}^{\dagger}$  ${}^{\dagger}$
BERT (Devlin et al., ${}^{\spadesuit}$ 81.6 81.7 78.4  ${}^{\dagger}$  ${}^{\dagger}$
BERTweet (Nguyen et al., 2020) 82.4 81.0 78.1  ${}^{\dagger}$  ${}^{\dagger}$ 69.2 70.7 69.0 67.9
JointCL (Liang et al., ${}^{\heartsuit}$  ${}^{\dagger}$  ${}^{\dagger}$
@@
[…]
@@ 5.2 Filtering Strategies
(a) Number of Demonstrations
(b) Number of Experts
(c) Number of Discussion Turns

### v3

@@ 1 Introduction
We evaluate DEEM across both single-target and multi-target stance detection tasks on three widely used datasets, including P-Stance (Li et al., 2021), SemEval-2016 (Mohammad et al., 2016), and MTSD (Sobhani et al., 2017). Experimental results demonstrate that DEEM with dynamic experienced experts can gain substantial improvement across all datasets. Furthermore, it also outperforms reasoning with self-consistency that requires multiple responses and shows potential for reducing the bias of LLMs. Code is available at https://github.com/THUNLP-MT/DEEM.
@@ 3 Methods
Figure 3: The overall pipeline of our DEEM method. We first use the prompt and training sentences to generate diverse experts (§$\S$ 3.1). Then we filter experienced experts according to their occurrence numbers and performance (§$\S$ 3.2). Finally, we build the sentence-expert pairs and retrieve the experienced experts for each new sentence (§$\S$ 3.3).
@@ 3.1 Generating Diverse Experts
$\mathcal{E}_{j},\hat{l}_{j}={\mathcal{M}}({\rm prompt}\oplus s_{j}\oplus t_{j}),$ (2)
@@ 4.1 Experimental Setups
FT BiCond (Augenstein et al., 2016) 73.0 69.4 64.6  32.7  30.5 - - - -
BERT (Devlin et al., 2019) 81.6 81.7 78.4  49.6  40.1 - - - -
BERTweet (Nguyen et al., 2020) 82.4 81.0 78.1  50.9  42.2 69.2 70.7 69.0 67.9
JointCL (Liang et al., 2022b) - - -  54.8  50.5 - - - -
@@
Metric. Following Liang et al. (2022a) and Zhang et al. (2023a), we use the F1avg, i.e., the average of F1-score on the label ‘favor’ and ‘against’, as the metric for evaluation and comparison.
@@ 5.2 Filtering Strategies
(a) Number of Demonstrations (b) Number of Experts (c) Number of Discussion Turns